A car uses 9 gallons of gasoline on a trip of 108 miles. At that rate, how much gasoline will a trip of 192 miles require? Set up a proportion and solve.
Answer by Nate(3500)

9 gallons / 108 miles = x gallons / 192 miles
9 / 108 = x / 192
1 / 12 = x / 192
16 gallons = x